# QMambaBSR: Burst Image Super-Resolution with Query State Space Model

### v1

@@ Introduction
Table 1: Performance comparison of existing methods on Synthetic BurstSR dataset.
$h_{t}$ $=\sum_{j=0}^{t}\left[\prod_{i=j+1}^{t}\exp(\Delta_{\text{base}_{i}}A)\right]\tilde{f}(\Delta_{\text{base}_{j}})\tilde{g}(B_{\text{base}_{j}})x_{\text{cur}_{j}},$ (5)
$y_{t}$ $={C\sum_{j=0}^{t}\left[\prod_{i=j+1}^{t}\exp(\Delta_{\text{base}_{i}}A)\right]\tilde{f}(\Delta_{\text{base}_{j}})\tilde{g}(B_{\text{base}_{j}})x_{\text{cur}_{j}}}$ (6)
$\quad+{Dx_{\text{cur}_{t}}},$
where $\tilde{f}$ and $\tilde{g}$ represent the functions of $\Delta$ and $B$ corresponding to the zero-order hold of $B$, as follows:
$\tilde{f}(\Delta_{\text{base}_{t}})$ $=(\Delta_{\text{base}_{t}}A)^{-1}\left(\exp(\Delta_{\text{base}_{t}}A)-I\right)\Delta_{\text{base}_{t}},$ (7)
$\tilde{g}(B_{\text{base}_{t}})$ $=B_{\text{base}_{t}}.$
[…]
We modify the RSSB block  (Guo et al. 2024) by integrating our proposed QSSM with four scanning directions and using channel attention to enhance channel interaction.
@@
Table 2: Performance comparison of existing methods on RealBSR-RGB and RealBSR-RAW datasets.
Multi-Scale Fusion Module
Considering the presence of sub-pixel information across various scales within the intricate details of images, we propose a novel Multi-scale Fusion Module (MSFM), as shown in Fig. 2 (c). This module is designed to effectively integrate multi-scale sub-pixel information from the current frames, thereby enhancing the capability for detailed image reconstruction. The MSFM comprises three distinct branches: a Convolutional Neural Network (CNN), a State Space Model (SSM) with diverse scanning orientations, and a channel Transformer. To begin with, a $3\times 3$ convolution is utilized for the fusion of local sub-pixel features. The SSM is introduced to efficiently learn and integrate sub-pixel features along both horizontal and vertical axes. Furthermore, considering the attenuation characteristics of the A matrix within the SSM when dealing with long-range perception, we concurrently employ a Transformer block to augment the network’s proficiency in capturing global information. The mathematical formulation of the MSFM is as follows:
where $w$ represents the balancing factors.
@@
Adaptive Up-sampling Module
After the aforementioned processes, the sub-pixel structural information from burst LR images is extracted and distributed in the feature space. The next critical challenge is to utilize this valuable sub-pixel information to adaptively upsample the image resolution and incorporate sub-pixel details into the high-resolution image. Existing state-of-the-art methods, such as Burstormer (Dudhane et al. 2023) and BIPNet (Dudhane et al. 2022), simply employ interpolation, transposed convolution, or pixel shuffle techniques for resolution upsampling. However, these methods lack the capability to perceive the distribution of sub-pixels in the feature space, leading to an inability to adaptively reconstruct fine details. Therefore, we introduce a novel Adaptive Up-sampling (AdaUp) module that perceives the distribution of sub-pixels in the spatial domain and adaptively adjusts the up-sampling kernel, thereby achieving higher-quality image detail, as shown in Fig. 2 (b). Specifically, we first adaptively perceive the distribution of sub-pixels $L\in\mathbb{R}^{B\times C_{\text{in}}\times 1\times 1}$ from the input features $X\in\mathbb{R}^{B\times C_{\text{in}}\times H\times W}$ by adaptive pooling. We then perform sequence feature interaction on $L$ to obtain the output channel feature distribution sequence $L_{1}\in\mathbb{R}^{B\times C_{\text{out}}\times 1\times 1}$. Subsequently, we apply both the input distribution sequence and the output distribution sequence to the upsampling transposed convolution kernel $W\in\mathbb{R}^{B\times C_{\text{in}}\times C_{\text{out}}\times 3\times 3}$ using broadcasting, thereby endowing the kernel with feature perception capability. Finally, we obtain the high-resolution output through the upsampling transposed convolution kernel. The corresponding formulas are as follows:
$L=\text{AdaptivePooling}(X)$ (9)
$L_{1}=\text{Conv}_{1\times 1}(L)$ (10)
$W_{f}=(W\odot L)\odot L_{1}$ (11)
$y=\text{Trans-Conv}(W_{f},X)$ (12)
where $\odot$ represents element-wise multiplication. Thus, AdaUp can leverage the underlying content information from input frames at different channels and utilize it to get better performance than the mainstream up-sampling operations, pixel shuffle, or interpolations (Carlson and Fritsch 1985; Schaefer, McPhail, and Warren 2006).
Experiments and Analysis
Experimental Settings
@@
Implementation details. We evaluate the effectiveness of our proposed method on four public burst image super-resolution benchmarks, encompassing both synthetic and real datasets: synthetic BurstSR (Bhat et al. 2021a), Real BurstSR (Bhat et al. 2021a), RealBSR-RAW (Wei et al. 2023), and RealBSR-RGB  (Wei et al. 2023). To ensure fairness, we follow (Dudhane et al. 2023)for training and evaluation. More details of datasets and data processing can be found in Appendix A section. Following (Dudhane et al. 2023), we train the model from scratch on the synthetic Burst SR dataset for 300 epochs, using the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We employ a cosine annealing strategy to gradually decrease the learning rate from $3\times 10^{-4}$ to $10^{-6}$ and set the training patch size to $48\times 48$. For the Real Burst SR dataset, we follow (Bhat et al. 2021a) to fine-tune the model pre-trained on synthetic Burst SR for 60 epochs, maintaining the same training setting as the synthetic Burst SR but adjusting the learning rate to $1\times 10^{-6}$ and the training patch size to $56\times 56$. For the RealBSR-RAW and RealBSR-RGB datasets, we follow (Wei et al. 2023) to train from scratch for 100 epochs, using the same training setting as the synthetic Burst SR, with a training patch size of $80\times 80$. We set the batch size to 8, and the burst size to 14, and all experiments are conducted on 8 V100 GPUs.
@@
Metric. Following previous works (Bhat et al. 2021a; Wei et al. 2023), we use reference metrics to evaluate performance, including PSNR, SSIM, and LPIPS.
@@
Compared methods. To comprehensively demonstrate the superiority of our proposed method, we compare our QMambaBSR with ten classic and state-of-the-art (SOTA) BurstSR approaches HighRes-net (Deudon et al. 2020), DBSR (Bhat et al. 2021a), LKR (Lecouat, Ponce, and Mairal 2021), MFIR (Bhat et al. 2021b), BIPNet (Dudhane et al. 2022), AFCNet (Mehta et al. 2022), FBAnet (Wei et al. 2023), Burstormer (Dudhane et al. 2023), RBSR (Wu et al. 2023), GMTNet (Mehta et al. 2023).
Quantitative and Qualitative Results
@@
Results on the Synthetic BurstSR dataset. As shown in Table 1, our method outperforms existing BurstSR methods, achieving the best performance. For example, compared to the existing SOTA method, Burstormer, our method achieves a 0.29 dB improvement in PSNR. Furthermore, to further demonstrate the visual superiority of our method, we present a visual comparison with existing methods in Fig. 3. We can observe that the RAW low-resolution images exhibit significant noise and severe detail loss, as illustrated in the window area of Fig. 3. Compared to existing methods, our method demonstrates superior performance in reconstructing textures and details in the window area. Additionally, in the green stripes region at the bottom of Fig. 3, the substantial noise in the base frame leads existing methods to leave artifacts. However, our method more effectively distinguishes between noise and sub-pixels, resulting in high-resolution images rich in details and free of artifacts, thereby demonstrating the visual superiority of our method.
@@
Results on RealBSR-RGB and RealBSR-RAW. As shown in Table 2, our method consistently outperforms existing methods on these two real benchmarks, achieving the best performance. For RealBSR-RAW, our method surpasses FBANet and Burstormer in PSNR and linear-PSNR by 0.268 dB and 0.258 dB, respectively. For RealBSR-RGB, our method surpasses FBANet and Burstormer in PSNR by 0.204 dB. Furthermore, as shown in Fig. 4, our method demonstrates superior performance in detail reconstruction and artifact suppression. This validates the effectiveness of our method in real-world scenarios, highlighting its superiority and practicality. More results on Real BurstSR and qualitative results will be presented in the appendix.
@@
Table 3: Ablation experiment on proposed modules.
Ablation Study
@@
Fusion Concat 39.87
PBFF (Dudhane et al. 2022) 40.77
NRFE (Dudhane et al. 2023) 41.89
Ours 42.44
Up-sampler Pixelshuffle 42.22
Transposed conv 42.19
Ours 42.44
Table 4: Comparison with existing modules.
Comparison with Existing Modules. To verify the effectiveness of proposed module, we replaced it with existing fusion and up-sampling modules. As shown in Table 4, in the fusion stage, compared to PBFF (Dudhane et al. 2022) or NRFE (Dudhane et al. 2023), our QSSM and MSFM modules are able to better exploit the inter-frame consistency of sub-pixel distribution while denoising, resulting in PSNR improvements of 1.67 dB and 0.55 dB, respectively. In contrast to static upsampling like pixel shuffle and transposed convolution, our AdaUp module enhances the network’s ability to perceive scene-specific sub-pixel distributions, leading to a PSNR improvement of 0.22 dB.
Figure 5: User study of reconstructed real HR images.
User study
To demonstrate the superiority of our proposed method in reconstructing visually pleasing images, we conduct a user study involving 10 real burst images from existing real benchmarks. Twenty volunteers rate the similarity and quality between each reconstructed image and the ground truth (GT) on a scale from 0 (visually unsatisfactory, completely dissimilar) to 10 (visually satisfactory, very similar). We then aggregate the scores from all volunteers, and the results are shown in Figure 5. Compared to existing methods such as Burstormer and BIPNet, our proposed method adaptively extracts sub-pixels to achieve the best visual effects, obtaining the best average score of 8.56.
Conclusion

### v2

@@ 4 Experiments and Analysis
Table 1: Performance comparison of existing methods on Synthetic BurstSR dataset for ×4 BurstSR.
@@
Table 2: Performance comparison of existing methods on RealBSR-RGB and RealBSR-RAW datasets for ×4 BurstSR.
4.1 Experimental Settings
@@
Implementation details. We evaluate the effectiveness of our proposed method on four public burst image super-resolution benchmarks, encompassing both synthetic and real datasets: synthetic BurstSR [1], Real BurstSR [1], RealBSR-RAW [45], and RealBSR-RGB  [45]. To ensure fairness, we follow [12] for training and evaluation. More details of datasets and data processing can be found in Appendix A section. Following [12], we train the model from scratch on the synthetic BurstSR dataset for 300 epochs, using the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We employ a cosine annealing strategy to gradually decrease the learning rate from $3\times 10^{-4}$ to $10^{-6}$ and set the training patch size to $48\times 48$. For the Real BurstSR dataset, we follow [1] to fine-tune the model pre-trained on synthetic BurstSR for 60 epochs, maintaining the same training setting as the synthetic BurstSR but adjusting the learning rate to $1\times 10^{-6}$ and the training patch size to $56\times 56$. For the RealBSR-RAW and RealBSR-RGB datasets, we follow [45] to train from scratch for 100 epochs, using the same training setting as the synthetic BurstSR, with a training patch size of $80\times 80$. We set the batch size to 8, and the burst size to 14, and all experiments are conducted on 8 V100 GPUs.
@@
Metric. Following previous works [1, 45], we use reference metrics to evaluate performance, including PSNR and SSIM.
@@
Compared methods. To comprehensively demonstrate the superiority of our proposed method, we compare our QMambaBSR with ten classic and state-of-the-art (SOTA) BurstSR approaches HighRes-net [8], DBSR [1], LKR [21], MFIR [2], BIPNet [11], AFCNet [30], FBAnet [45], Burstormer [12], RBSR [46], GMTNet [31].
4.2 Quantitative and Qualitative Results
@@
Results on the Synthetic BurstSR dataset. As shown in Table 1, our method outperforms existing BurstSR methods, achieving the best performance. For example, compared to the existing SOTA method, Burstormer, our method achieves a 0.29 dB improvement in PSNR. Furthermore, to further demonstrate the visual superiority of our method, we present a visual comparison with existing methods in Figure 4. We can observe that the RAW low-resolution images exhibit significant noise and severe detail loss, as illustrated in the window area of Figure 4. Compared to existing methods, our method demonstrates superior performance in reconstructing textures and details in the window area. Additionally, in the green stripes region at the bottom of Figure 4, the substantial noise in the base frame leads existing methods to leave artifacts. However, our method more effectively distinguishes between noise and sub-pixels, producing more detailed, artifact-free high-resolution images, thereby demonstrating the visual superiority of our method.
@@
Results on RealBSR-RGB and RealBSR-RAW. As shown in Table 2, our method consistently outperforms existing methods on these two real benchmarks, achieving the best performance. For RealBSR-RAW, our method surpasses FBANet and Burstormer in PSNR and linear-PSNR by 0.268 dB and 0.258 dB, respectively. For RealBSR-RGB, our method surpasses FBANet and Burstormer in PSNR by 0.204 dB. Furthermore, as shown in Figure 5, the results demonstrate the superior performance of our method in detail reconstruction and artifact suppression. This validates the effectiveness of our method in real-world scenarios, highlighting its superiority and practicality. More results on Real BurstSR and qualitative results will be presented in the appendix.
4.3 Ablation Study
Comparison with Existing Modules. To verify the effectiveness of proposed module, we replaced it with existing fusion and up-sampling modules. To effectively represent the performance of various methods while also considering training time, we train for 100 epochs on the synthetic dataset. As shown in Table 4, in the fusion stage, compared to PBFF [11] or NRFE [12], our QSSM and MSFM modules are able to better exploit the inter-frame consistency of sub-pixel distribution while denoising, resulting in PSNR improvements of 1.56 dB and 0.41 dB, respectively. In contrast to static upsampling like pixel shuffle and transposed convolution, our AdaUp module enhances the network’s ability to perceive scene-specific sub-pixel distributions, leading to a PSNR improvement of 0.16 dB.
@@ 4.3 Ablation Study
Table 3: Ablation experiment on proposed core modules.
Evaluation about MSFM. To validate the effectiveness of different branches within the proposed MSFM, we conducted ablation experiments with various internal branch setting. The experimental results are shown in Table 5. To minimize training time while ensuring experimental validity, the ablation experiments were conducted on a synthetic dataset and trained for 100 epochs. As shown in the table, we use only the convolution in the MSFM module as the baseline. Adding the transformer module resulted in a PSNR increase of 0.39 dB. When the transformer was replaced with SSM, the PSNR improved by 0.44 dB compared to using only the conv method. Under our proposed MSFM, the PSNR increased by 0.56 dB, clearly demonstrating the effectiveness of both the SSM and transformer branches in the MSFM.
Figure 6: User study of reconstructed real HR images.
@@
Fusion Concat 39.85 0.93
PBFF [11] 40.57 0.94
NRFE [12] 41.72 0.96
Ours 42.13 0.96
Up-sampler Pixelshuffle 41.89 0.96
Transposed conv 41.97 0.96
Ours 42.13 0.96
Table 4: Comparison with existing modules on fusion and upsampling process stage.
4.4 User study
To effectively illustrate the superiority of our proposed method in reconstructing visually pleasing images, we conducted a user study involving 10 real burst images selected from well-established benchmarks. Twenty volunteers participated in this study, tasked with rating the similarity and quality between each reconstructed image and the ground truth (GT). They used a detailed scale ranging from 0, indicating visually unsatisfactory and completely dissimilar images, to 10, representing visually satisfactory and highly similar images. We then meticulously aggregated the scores from all volunteers, and the results are depicted in Figure 6. When compared to existing methods such as Burstormer and BIPNet, our proposed method stands out by adaptively extracting sub-pixels to achieve superior visual effects. This approach earned our method the highest average score of 8.56, underscoring its effectiveness.
[…]
× × 41.57 0.95
× 41.96 0.96
× 42.01 0.96
42.13 0.96
Table 5: Ablation studies on our proposed MSFM.
5 Conclusion
@@
In future work, we plan to utilize state space models further to enhance the alignment stage in burst image super-resolution and explore more efficient and coherent integration of our architecture across various vision tasks. Additionally, we aim to unlock our method to other burst image restoration tasks, such as denoising, HDR, and more.